OpenStudy (anonymous):

how do you find the point of intersection y=x^2 , y=x^3

OpenStudy (anonymous):

set the two equations for y as equal to each other

OpenStudy (anonymous):

x^2 =x^3 and then (x^2)^1/3 ?

OpenStudy (anonymous):

no then \[x^3-x^2=0\] \[x^2(x-1)=0\] etc